Weaver Company’s predetermined overhead rate is $15.00 per direct labour-hour, and its direct labour wage rate is $10.00 per hour. The following information pertains to Job A-200:

    
  Direct materials $315
  Direct labour $210

Required:

1. What is the total manufacturing cost assigned to Job A-200?

2. If Job A-200 consists of 40 units, what is the average cost assigned to each unit included in the job?

Score answer Visit question map

Question 1 of 2 Total
Answer #1

1. 840

2. 21

A DU - Total manufacturing cost = Direct Materials 315 + Direct Labor 210 + Overhead (21*15) = 315+210+315 = 840 Cost per uni